In the recent years, PT Angus has purchased three delivery trucks Because of frequent employee turnover in the accounting department, a different accountant was in charge of selecting the depreciation method of each delivery truck, and various methods have been used. Information concerning the delivery trucks is summarized in the table below Delivery Acquired Cost Salvage Useful Life Depreciation Method Truck Value (in years) A ly 1, 2012 $125,000S 10,000 B January 1,2013 S 95,000S 7,500 10 Straight-line Double declining- balance C January 1, 2013S 96,000 S 6,000 Units -of-activity For the delivery truck number C, the truck is expected to be driven 150,000 miles. Actual miles of use in the first 3 years were: 2013: 2,800, 2014: 3,200; 2015: 3,300 Required: 1. Prepare journal entries for purchasing those delivery trucks in cash 2. Compute the amount of depreciation expense of each delivery truck per December 31, 2013
Answer #1

Requirement 1

Date

General Journal

Debit

Credit

July 1 2012

Delivery truck-A

$ 1,25,000.00

Cash

$   1,25,000.00

(Delivery truck purchased)

January 1 2013

Delivery truck-B

$     95,000.00

Delivery truck-C

$     96,000.00

Cash

$   1,91,000.00

(Delivery truck purchased)

Requirement 2

2013

Delivery truck A

$ 11500

Delivery truck B

$ 38000

Delivery truck C

$ 1680

Straight line Method

A

Cost

$   1,25,000.00

B

Residual Value

$      10,000.00

C=A - B

Depreciable base

$   1,15,000.00

D

Life [in years left ]

                        10

E=C/D

Annual SLM depreciation

$      11,500.00

Double declining Method

A

Cost

$      95,000.00

B

Residual Value

$         7,500.00

C=A - B

Depreciable base

$      87,500.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$      17,500.00

F=E/C

SLM Rate

20.00%

G=F x 2

DDB Rate

40.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

2013

$ 95,000.00

40.00%

$ 38,000.00

$ 57,000.00

$ 38,000.00

Units of Production method

A

Cost

$    96,000.00

B

Residual Value

$      6,000.00

C=A - B

Depreciable base

$    90,000.00

D

Usage in units(in Miles)

150000

E

Depreciation per Mile

$ 0.60

Year

Book Value

Usage

Depreciation expense

Ending Book Value

Accumulated Depreciation

2017

$ 96,000.00

2800

$ 1,680.00

$ 94,320.00

$ 1,680.00